# Optimizing Diffusion Noise Can Serve As Universal Motion Priors

### v1

@@ 1 Introduction
Many applications of great interest to the motion modeling community can be framed as “finding a plausible motion that fulfills a set of criteria”, typically formulated as minimizing a cost function addressing the given constraints. These include but are not limited to generating motions that follow a trajectory, target locations for joints, keyframes, or avoid obstacles; denoising, completing missing parts of a motion or editing an existing one. Though they may seem diverse, a unified framework should be able to address these diverse yet relevant tasks, which typically follow a task-agnostic motion prior that ensures plausible motions and task-specific cost functions. Such high-quality motion priors are highly sought after, often with efforts that focus on improving architectures. Yet, a truly versatile framework utilizing the underlying motion prior expressively is still lacking.
@@
Among many human motion priors, the diffusion-based models [37, 60, 42, 29, 73, 12] have become the most prominent ones by achieving impressive performance on certain motion modeling tasks. Avatar Grow Legs [12] and EgoEgo [37] can generate full-body motions from limited joint specifications, such as head and hand poses, but their specific scope precludes their use as general motion priors. PriorMDM [60] introduces a more flexible setting; however, the underlying root-relative motion representation requires dense condition signals for individual joints or keyframes, limiting its use to a small set of applications. OmniControl [73] addresses this limitation by augmenting the motion diffusion model with another network to encode input conditions. While this improves the fidelity and quantity of control signals, it is not always straightforward to support many other constraints. For instance, generating locomotion in a scene with obstacles requires a perception of the physical scene, which is unclear how to cover arbitrary scenes effectively. It is only exacerbated when the scene dynamically changes over time, such as a scene populated by other human agents. GMD [29] presents a more versatile approach by limiting the conditioning to task-based objectives, which theoretically supports arbitrary use cases. Yet there is a trade-off between this flexibility and the fidelity of controls and quality of motions.
@@
To this end, we propose a simple yet effective approach to utilize a motion diffusion model as a motion prior. By treating the denoising process as a black box, we can frame motion-related tasks, such as motion editing and refinement, as optimization problems on the latent manifold of the diffusion model, similar to other classes of generative models such as GANs [46] and VAEs [55] where optimization is seen as iteratively updating the solution on the learned motion manifold (Fig. 2a). In this work, we demonstrate that it is possible and feasible to back-propagate gradients through the full-chain diffusion process, and then optimize the noise vector based on user-provided criteria in the motion space. This simple approach, dubbed Diffusion Noise Optimization (DNO), which we show to be effective at editing and preserving contents, is a unified and versatile framework that supports a wide variety of applications without the need for fine-tuning the underlying motion model for each specific application. By changing the optimization objective, defined as any differentiable loss function computed on the output motion, DNO enables the diffusion model to effectively serve as a motion prior.
@@ 2 Related Works
Motion synthesis, editing, and completion. The human motion generation task aims to generate motions either conditionally or unconditionally [47, 74, 81, 82]. Various conditioning signals have been explored such as partial poses [13, 18], trajectories [70, 80, 30, 60], images [55, 7], music [34, 35, 38], objects [72], action labels [48, 14], or text [1, 16, 17, 31, 49]. Recently, the focus has been on generating motion based on natural-language descriptions using diffusion-based models [66, 5]. These models, utilizing the CLIP model [53], have shown significant improvements in text-to-motion generation [6, 76, 79] and support conditioning on partial motions or music [67, 2]. Alternatively, the motion can be treated as a new language and embedded into the language model framework [77, 26] Nevertheless, they lack the ability to handle spatial conditioning signals, such as keyframe locations or trajectories. GMD [29] handles this problem with classifier-based guidance at inference time to steer the denoising process toward the target conditions. OmniControl [73] combined GMD with ControlNet [78] to improve realism but limits the conditioning signals to text and partial motion observations. However, motion editing under the motion synthesis framework remains unexplored due to the lack of an explicit mechanism to retain the original motion content.
@@
In parallel, the generation of full-body poses from sparse tracking signals of body joints has gained considerable interest within the community. Previous work such as EgoEgo [37], AGRoL [12], and AvatarPoser [27] while showing impressive results, tend to be specialized models trained explicitly for the motion completion or denoising task. Therefore, it is unclear how we can leverage the motion priors in these trained models for solving other tasks. In this work, we focus on using the trained motion diffusion model to tackle various motion-related tasks under a unified framework including editing, completion, and refinement.
@@
Diffusion models and guidance. Diffusion-based probabilistic generative models (DPMs), a class of generative models based on learning to progressively denoising the input data, [22, 61, 63, 64] have gained significant attention across multiple fields of research. They have been successfully applied to tasks such as image generation [11], image super-resolution [57, 36], speech synthesis [33, 33, 51], video generation [24, 23], 3D shape generation [50, 71], and reinforcement learning [25]. The growing interest in the diffusion-based model stems from their superior results and impressive generation controllability, for example, in text-conditioned generation [54, 56, 58] and image-conditioned editing [44, 8, 4, 20, 3]. In terms of conditioning, there are various methods for the diffusion-based models such as imputation and inpainting [9, 8, 44], classifier-based guidance [11, 9], and classifier-free guidance [56, 58, 54, 21]. For refinement, SDEdit [44] enables the repetition of the denoising process to gradually improve output quality but lacks the ability to provide editing guidance. Recently, DOODL [68] demonstrates a direct latent optimization approach for image editing with the help of an invertible ODE [69]. Being inspired by that work, we propose a related method with an improved optimization algorithm that speeds up optimization. DNO can effectively be used in various motion-related tasks making it suitable as a versatile human motion priors. Additionally, we discover that an invertible ODE is not required which makes DNO simpler and convenient to use with minimal effort.
@@ 3.1 Motion generation with diffusion model
While diffusion models are stochastic, there exist deterministic sampling processes that share the same marginal distribution. These processes include those defined by probability flow ODE [64] or by reformulating the diffusion process to be non-Markovian as in DDIM [62]. To obtain a sample based on deterministic sampling, we can solve the associated ODE using an ODE solver starting from ${\mathbf{x}_{T}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@ 3.2 Diffusion model inversion
Diffusion inversion is a process that retrieves the corresponding noise map ${\mathbf{x}_{T}}$ given an input ${\mathbf{x}_{0}}$. Under deterministic sampling, the associated ODE not only describes the denoising process from ${\mathbf{x}_{T}}$ to ${\mathbf{x}_{0}}$ but also its reverse from ${\mathbf{x}_{0}}$ to ${\mathbf{x}_{T}}$, which is achievable by solving the ODE backward. However, the fidelity of this inversion relies on the smoothness assumption $d_{\theta}({\mathbf{x}_{t}})\approx d_{\theta}({\mathbf{x}_{t-1}})$, which is unlikely to hold true when solving the ODEs with just a few discretization steps. For tasks requiring multiple back-and-forth evaluations between ${\mathbf{x}_{T}}$ and ${\mathbf{x}_{0}}$, an alternative inversion method is available [69].
@@ 3.3 Motion representation
The relative-root representation [15] has been widely adopted for text-to-motion diffusion models [66]. The representation is a matrix of human joint features over the motion frames with shape $D\times M$, where $D=263$ and $M$ are the representation size and the number of motion frames, respectively. Each motion frame represents root relative rotation and velocity, root height, joint locations, velocities, rotations, and foot contact labels. As the relative representation abstracts away the absolute root location, it can improve the generalization of motion models but makes the controllable generation more challenging [66, 29].
@@ 4 Diffusion Noise Optimization
Such optimization has led to the success in image editing of GANs [28, 19], enabled by its smooth latent space $\mathbf{z}$ whose mapping is parameterized by a powerful generative model $f(\mathbf{z})$ and learned from a large dataset. In this work, we show that the latent optimization extends well to $f$, parameterized as pretrained diffusion models in the motion domain.
For diffusion models, our choice of the latent variable is the diffusion’s noise at time $T$, $\mathbf{z}={\mathbf{x}_{T}}$. While ${\mathbf{x}_{T}}$ is not perfectly smooth [52, 62], ${\mathbf{x}_{T}}$ offers the highest-level abstraction readily available. Now, getting the output $\mathbf{x}$ from a diffusion model $d(\cdot)$ requires solving an ordinary differential equation (ODE) using an ODE solver [64]. We now see Equation 1 as
@@
This allows us to approach many motion tasks simply by adjusting the task-specific criterion $\mathcal{L}$ while keeping the motion model, $d$, intact. In this work, we use the DDIM-ODE [62] and its solver.
This optimization is iteratively solved using gradient descent. Starting from a certain noise ${\mathbf{x}_{T}}$, we solve the ODE, arrive at a prediction $\mathbf{x}$, and evaluate the criterion function $\mathcal{L}(\mathbf{x})$. Then, we obtain the gradient $\nabla_{{\mathbf{x}_{T}}}\mathcal{L}(\mathbf{x})$ by backpropagating through the ODE solver. An optimizer updates ${\mathbf{x}_{T}}$ based on the gradient, possibly with a small random perturbation to encourage exploration [28, 68]. We repeat this until convergence. The final output is the motion obtained by solving the ODE one last time with the optimized ${\mathbf{x}_{T}}$. We call the above algorithm Diffusion Noise Optimization (DNO) and summarize it in Algo. 1.
Maintaining the intermediate activations for solving the ODE during backpropagation can be memory-intensive. This issue can be addressed with an invertible ODE [69, 68], though at the cost of significant model complexity. In the motion domain, impressive results can be achieved from an ODE solver with a minimal number of function evaluations and feasible memory overhead. Given the ongoing efforts to reduce diffusion sampling steps, as evidenced in distillation [59, 65, 39] and high-order solvers studies [41, 40, 83], our simple design, despite requiring back-propagation through all steps, can become increasingly relevant and applicable.
Optimization through ODE solver. Empirically, gradients via backpropagation through the ODE solver have norms spanning multiple orders of magnitude, making the optimization unstable and slow. We instead propose to normalize the gradient to have a unit norm, which optimizes faster in practice. Normalized gradient methods are also found to help escape saddle-point in the loss landscape faster [45, 10]. When combined with momentum, the update becomes a moving average of directions, which is robust to gradient norm outliers.
@@ 5.2 Motion refinement and completion
Since this optimization begins with a random noise ${\mathbf{x}_{T}}$, the initial prediction can be far from the desired predicted motion, requiring significant changes on ${\mathbf{x}_{T}}$ with several optimization steps, which tend to correlate neighboring noises and reduce the representation capacity. We empirically find that regularizing the noise to decorrelate each latent dimension alleviates the foot skating problem. Inspired by StyleGAN2 [28], we introduce a latent decorrelation loss across motion frames:
@@ 6 Experiments
Datasets. When applicable, we evaluate generated motions on the HumanML3D [16] dataset, which contains 44,970 motion annotations of 14,646 motion sequences from AMASS [43] and HumanAct12 [14] datasets.
Motion diffusion model. As DNO is a plug-and-play method that is model-agnostic, it can potentially work with any trained motion diffusion model. We choose the Motion Diffusion Model (MDM) [66], which is trained on the HumanML3D dataset for our experiments. However, we found that the released MDM produced inconsistent output between checkpoints. Thus, we retrained MDM with Exponential Moving Averaging [22] to help stabilize the model, with no further modifications.
DNO implementation details. We use Adam optimizer [32] with a learning rate of $0.05$, a linear warm-up for the first $50$ steps, and a cosine learning rate schedule to zero for the entire optimization. We use a unit-normalized gradient. We set the coefficients $\lambda_{\text{cont}}=0.01,\lambda_{\text{decorr}}=10^{3},\lambda_{\text{obs}}=1.0$, and the perturbation amount $\gamma=0$. This design choice is explored in Section 6.3.
For editing tasks, we obtain the initial ${\mathbf{x}_{T}}$ from DDIM-100 inversion on MDM [66] without text conditions. The subsequent optimization is run for $300$ steps, which takes approximately $3$ minutes on an Nvidia 3090 GPU. For refinement tasks, the optimization is run for $500$ steps. DNO optimization with DDIM-10 and a batch size of $16$ requires $18$ GB of GPU memory. Optimization in all tasks is also done without text conditions to the diffusion model.

### v2

@@ 1 Introduction
Many applications of great interest to the motion modeling community can be framed as “finding a plausible motion that fulfills a set of criteria”, typically formulated as minimizing a cost function addressing the given constraints. These include but are not limited to generating motions that follow a trajectory, target locations for joints, keyframes, or avoiding obstacles; denoising, completing missing parts of a motion or editing an existing one. Though they may seem diverse, a unified framework should be able to address these diverse yet relevant tasks, which typically follow a task-agnostic motion prior that ensures plausible motions and task-specific cost functions. Such high-quality motion priors are highly sought after, often with efforts that focus on improving architectures. Yet, a truly versatile framework utilizing the underlying motion prior expressively is still lacking.
@@
Among many human motion priors, the diffusion-based models [Li2023-dn, Shafir2024-qh, ma2022mofusion, Karunratanakul2023-nb, Xie2024-zz, Du2023-kh] have become the most prominent ones by achieving impressive performance on certain motion modeling tasks. Avatar Grow Legs [Du2023-kh] and EgoEgo [Li2023-dn] can generate full-body motions from limited joint specifications, such as head and hand poses, but their specific scope precludes their use as general motion priors. PriorMDM [Shafir2024-qh] introduces a more flexible setting; however, the underlying root-relative motion representation requires dense condition signals for individual joints or keyframes, limiting its use to a small set of applications. OmniControl [Xie2024-zz] addresses this limitation by augmenting the motion diffusion model with another network to encode input conditions. While this improves the fidelity and quantity of control signals, it is not always straightforward to support many other constraints. For instance, generating locomotion in a scene with obstacles requires a perception of the physical scene, which is unclear how to cover arbitrary scenes effectively. It is only exacerbated when the scene dynamically changes over time, such as a scene populated by other human agents. GMD [Karunratanakul2023-nb] presents a more versatile approach by limiting the conditioning to task-based objectives, which theoretically supports arbitrary use cases. Yet there is a trade-off between this flexibility and the fidelity of controls and quality of motions.
@@
To this end, we propose a simple yet effective approach to utilize a motion diffusion model as a motion prior. By treating the denoising process as a black box, we can frame motion-related tasks, such as motion editing and refinement, as optimization problems on the latent manifold of the diffusion model, similar to other classes of generative models such as GANs [pan2023drag] and VAEs [Rempe2021-hp] where optimization is seen as iteratively updating the solution on the learned motion manifold (Fig. 2a). In this work, we demonstrate that it is possible and feasible to back-propagate gradients through the full-chain diffusion process, and then optimize the noise vector based on user-provided criteria in the motion space. This simple approach, dubbed Diffusion Noise Optimization (DNO), which we show to be effective at editing and preserving contents, is a unified and versatile framework that supports a wide variety of applications without the need for fine-tuning the underlying motion model for each specific application. By changing the optimization objective, defined as any differentiable loss function computed on the output motion, DNO enables the diffusion model to effectively serve as a motion prior.
@@ 2 Related Works
Motion synthesis, editing, and completion. The human motion generation task aims to generate motions either conditionally or unconditionally [safonova07motiongraphs, pavlakos2019expressive, yan2019convolutional, zhang2020perpetual, zhao2020bayesian]. Various conditioning signals have been explored such as partial poses [duan2021single, harvey2020robust], trajectories [wang2021synthesizing, zhang2021learning, kaufmann2020convolutional, Shafir2024-qh], images [Rempe2021-hp, chen2022learning], music [lee2019dancing, li2022danceformer, li2021ai], objects [wu2022saga], action labels [petrovich2021action, guo2020action2motion], scene [huang2023diffusion], or text [ahuja2019language2pose, guo2022t2m, guo2022tm2t, kim2022flame, petrovich2022temos]. Recently, the focus has been on generating motion based on natural-language descriptions using diffusion-based models [Tevet2023-ih, Chen2023-id]. These models, utilizing the CLIP model [radford2021learning], have shown significant improvements in text-to-motion generation [chen2022mld, yuan2022physdiff, zhang2022motiondiffuse] and support conditioning on partial motions or music [tseng2022edge, alexanderson2022listen]. Alternatively, the motion can be treated as a new language and embedded into the language model framework [zhang2023t2mgpt, jiang2023motiongpt] Nevertheless, they lack the ability to handle spatial conditioning signals, such as keyframe locations or trajectories. GMD [Karunratanakul2023-nb] handles this problem with classifier-based guidance at inference time to steer the denoising process toward the target conditions. OmniControl [Xie2024-zz] combined GMD with ControlNet [zhang2023controlnet] to improve realism but limits the conditioning signals to text and partial motion observations. However, motion editing under the motion synthesis framework remains unexplored due to the lack of an explicit mechanism to retain the original motion content.
@@
In parallel, the generation of full-body poses from sparse tracking signals of body joints has gained considerable interest within the community. Previous work such as EgoEgo [Li2023-dn], AGRoL [Du2023-kh], and AvatarPoser [jiang2022avatarposer] while showing impressive results, tend to be specialized models trained explicitly for the motion completion or denoising task. Therefore, it is unclear how we can leverage the motion priors in these trained models for solving other tasks. In this work, we focus on using the trained motion diffusion model to tackle various motion-related tasks under a unified framework including editing, completion, and refinement.
@@
Diffusion models and guidance. Diffusion-based probabilistic generative models (DPMs), a class of generative models based on learning to progressively denoising the input data, [Ho2020-ew, Sohl-Dickstein2015-hp, Song2019-xr, Song2021-yj] have gained significant attention across multiple fields of research. They have been successfully applied to tasks such as image generation [Dhariwal2021-bt], image super-resolution [Saharia2021-lm, Li2022-qr], speech synthesis [Kong2021-sn, Popov2021-qe], video generation [Ho2022-qm, Ho2022-gh], 3D shape generation [Poole2022-kx, Watson2022-vm], and reinforcement learning [Janner2022-no]. The growing interest in the diffusion-based model stems from their superior results and impressive generation controllability, for example, in text-conditioned generation [Ramesh2022-id, Rombach2022-nu, Saharia2022-ns] and image-conditioned editing [Meng2022-bn, Choi2021-oe, Brooks2023-wx, Hertz2022-td, Balaji2022-gh]. In terms of conditioning, there are various methods for the diffusion-based models such as imputation and inpainting [Chung2022-tm, Choi2021-oe, Meng2022-bn], classifier-based guidance [Dhariwal2021-bt, Chung2022-tm], and classifier-free guidance [Rombach2022-nu, Saharia2022-ns, Ramesh2022-id, ho2021classifierfree]. For refinement, SDEdit [Meng2022-bn] enables the repetition of the denoising process to gradually improve output quality but lacks the ability to provide editing guidance. Recently, DOODL [Wallace2023-ng] demonstrates a direct latent optimization approach for image editing with the help of an invertible ODE [Wallace2023-ja]. Being inspired by that work, we propose a related method with an improved optimization algorithm that speeds up optimization. DNO can effectively be used in various motion-related tasks making it suitable as a versatile human motion priors. Additionally, we discover that an invertible ODE is not required which makes DNO simpler and convenient to use with minimal effort.
@@ 3.1 Motion generation with diffusion model
While diffusion models are stochastic, there exist deterministic sampling processes that share the same marginal distribution. These processes include those defined by probability flow ODE [Song2021-yj] or by reformulating the diffusion process to be non-Markovian as in DDIM [Song2021-ac]. To obtain a sample based on deterministic sampling, we can solve the associated ODE using an ODE solver starting from ${\mathbf{x}_{T}}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$.
@@ 3.2 Diffusion model inversion
Diffusion inversion is a process that retrieves the corresponding noise map ${\mathbf{x}_{T}}$ given an input ${\mathbf{x}_{0}}$. Under deterministic sampling, the associated ODE not only describes the denoising process from ${\mathbf{x}_{T}}$ to ${\mathbf{x}_{0}}$ but also its reverse from ${\mathbf{x}_{0}}$ to ${\mathbf{x}_{T}}$, which is achievable by solving the ODE backward. However, the fidelity of this inversion relies on the smoothness assumption $d_{\theta}({\mathbf{x}_{t}})\approx d_{\theta}({\mathbf{x}_{t-1}})$, which is unlikely to hold true when solving the ODEs with just a few discretization steps. For tasks requiring multiple back-and-forth evaluations between ${\mathbf{x}_{T}}$ and ${\mathbf{x}_{0}}$, an alternative inversion method is available [Wallace2023-ja].
@@ 3.3 Motion representation
The relative-root representation [Guo2022-vx] has been widely adopted for text-to-motion diffusion models [Tevet2023-ih]. The representation is a matrix of human joint features over the motion frames with shape $D\times M$, where $D=263$ and $M$ are the representation size and the number of motion frames, respectively. Each motion frame represents root relative rotation and velocity, root height, joint locations, velocities, rotations, and foot contact labels. As the relative representation abstracts away the absolute root location, it can improve the generalization of motion models but makes the controllable generation more challenging [Tevet2023-ih, Karunratanakul2023-nb].
@@ 4 Diffusion Noise Optimization
Such optimization has led to the success in image editing of GANs [Karras2019-rv, He2019-lr], enabled by its smooth latent space $\mathbf{z}$ whose mapping is parameterized by a powerful generative model $f(\mathbf{z})$ and learned from a large dataset. In this work, we show that the latent optimization extends well to $f$, parameterized as pretrained diffusion models in the motion domain.
For diffusion models, our choice of the latent variable is the diffusion’s noise at time $T$, $\mathbf{z}={\mathbf{x}_{T}}$. While ${\mathbf{x}_{T}}$ is not perfectly smooth [Preechakul2022-ql, Song2021-ac], ${\mathbf{x}_{T}}$ offers the highest-level abstraction readily available. Now, getting the output $\mathbf{x}$ from a diffusion model $d(\cdot)$ requires solving an ordinary differential equation (ODE) using an ODE solver [Song2021-yj]. We now see Equation 1 as
@@
This allows us to approach many motion tasks simply by adjusting the task-specific criterion $\mathcal{L}$ while keeping the motion model, $d$, intact. In this work, we use the DDIM-ODE [Song2021-ac] and its solver.
This optimization is iteratively solved using gradient descent. Starting from a certain noise ${\mathbf{x}_{T}}$, we solve the ODE, arrive at a prediction $\mathbf{x}$, and evaluate the criterion function $\mathcal{L}(\mathbf{x})$. Then, we obtain the gradient $\nabla_{{\mathbf{x}_{T}}}\mathcal{L}(\mathbf{x})$ by backpropagating through the ODE solver. An optimizer updates ${\mathbf{x}_{T}}$ based on the gradient, possibly with a small random perturbation to encourage exploration [Karras2019-rv, Wallace2023-ng]. We repeat this until convergence. The final output is the motion obtained by solving the ODE one last time with the optimized ${\mathbf{x}_{T}}$. We call the above algorithm Diffusion Noise Optimization (DNO) and summarize it in Algo. 1.
Maintaining the intermediate activations for solving the ODE during backpropagation can be memory-intensive. This issue can be addressed with gradient checkpointing [Clark2024-bv] or an invertible ODE [Wallace2023-ja, Wallace2023-ng], at the cost of more computation or model complexity. In the motion domain, impressive results can be achieved from an ODE solver with a minimal number of function evaluations and feasible memory overhead. Given the ongoing efforts to reduce diffusion sampling steps, as evidenced in distillation [Salimans2022-um, Song2023-vx, Liu2024-ia] and high-order solvers studies [Lu2022-wv, Lu2022-rz, Zheng2023-lu], our simple design, despite requiring back-propagation through all steps, can become increasingly relevant and applicable.
Optimization through ODE solver. Empirically, gradients via backpropagation through the ODE solver have norms spanning multiple orders of magnitude, making the optimization unstable and slow. We instead propose to normalize the gradient to have a unit norm, which optimizes faster in practice. Normalized gradient methods are also found to help escape saddle-point in the loss landscape faster [Murray2019-cv, Cortes2006-lp]. When combined with momentum, the update becomes a moving average of directions, which is robust to gradient norm outliers.
@@
DNO fundamentally differs from a guided motion diffusion method [Karunratanakul2023-nb] in how the criterion $\mathcal{L}(\cdot)$ is computed at each iteration. For guided diffusion methods, the criterion is computed on an expected $\hat{x}=\mathbb{E}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ at a denoising step $t$ [song2023loss, yu2023freedom] that is $\mathcal{L}(\mathbf{x}_{t})=\mathbb{E}_{p(\mathbf{x}_{0}|\mathbf{x}_{t})}\mathcal{L}(\mathbf{x}_{0})\approx\mathcal{L}(\hat{\mathbf{x}})$, resulting in severe error when $\mathrm{Var}[\mathbf{x}_{0}|\mathbf{x}_{t}]$ is large. For DNO, the criterion is exactly computed on $\mathbf{x}$ after the full-chain denoising which eliminates the approximation error. We further discuss these differences in the supp. mat.
@@ 5.2 Motion refinement and completion
Since this optimization begins with a random noise ${\mathbf{x}_{T}}$, the initial prediction can be far from the desired predicted motion, requiring significant changes on ${\mathbf{x}_{T}}$ with several optimization steps, which tend to correlate neighboring noises and reduce the representation capacity. We empirically find that regularizing the noise to decorrelate each latent dimension alleviates the foot skating problem. Inspired by StyleGAN2 [Karras2019-rv], we introduce a latent decorrelation loss across motion frames:
@@ 6 Experiments
Datasets. When applicable, we evaluate generated motions on the HumanML3D [guo2022t2m] dataset, which contains 44,970 motion annotations of 14,646 motion sequences from AMASS [mahmood2019amass] and HumanAct12 [guo2020action2motion] datasets.
Motion diffusion model. As DNO is plug-and-play and model-agnostic, it works with any trained motion diffusion model. When paired with MDM [Tevet2023-ih], which is trained on the HumanML3D dataset, we name the combination DNO-MDM. We retrained MDM with Exponential Moving Averaging [Ho2020-ew] with no further modifications to help stabilize the model, which was previously found to produce inconsistent output between checkpoints.
DNO implementation details. We use Adam optimizer [kingma2014adam] with a learning rate of $0.05$, a linear warm-up for the first $50$ steps, and a cosine learning rate schedule to zero for the entire optimization. We use a unit-normalized gradient. We set the coefficients $\lambda_{\text{cont}}=0.01,\lambda_{\text{decorr}}=10^{3},\lambda_{\text{obs}}=1.0$, and the perturbation amount $\gamma=0$. This design choice is explored in Section LABEL:sec:ablation.
For editing tasks, we obtain the initial ${\mathbf{x}_{T}}$ from DDIM-100 inversion on MDM [Tevet2023-ih] without text conditions. The subsequent optimization is run for $300$ steps, which takes approximately $3$ minutes on an Nvidia 3090 GPU. For refinement tasks, the optimization is run for $500$ steps. DNO optimization with DDIM-10 and a batch size of $16$ requires $18$ GB of GPU memory. Optimization in all tasks is also done without text conditions to the diffusion model. Gradient checkpointing [Chen2016-vo, Clark2024-bv] can be used with DNO to reduce memory usage further for a low-memory GPU at the cost of more computation.